# ReCoRe: Regularized Contrastive Representation Learning of World Model

### v1

@@ 3.1.2 Intervention-invariant Regularizer
Model-based RL approaches such as DreamerV2, separate the training of the world model and controller. This makes model-based RL more sample efficient as compared to model-free RL, since future predictions from latent dynamics (rollouts) can be used to learn RL policy. However, the policy learning is disconnected from world-model learning. Therefore, the supervisory signal from policy does not optimize the encoder. While the image reconstruction loss in DreamerV2 caters for this issue for in-distribution learning. For out-of-distribution generalization, this creates an issue, since the policy was optimized for a different environment as compared to the world-model. Contrastive learning further elevates the issue, since it trains the encoder under different data augmentations while the world-model aims to predict the true image using latent dynamics. This leads to collapse of encoder features due to lack of correct supervisory signal, therefore the brute-force combination of world-model and contrastive loss fails as can be seen in Table 1, ReCoRe vs. ReCoRe-D).
To this end, we propose regularization in the form of intervention invariant auxiliary task to explicitly enforce invariance, which is robust against changes to the nuisance variables. For navigation tasks, we choose depth reconstruction $p_{\theta}(\widehat{d}_{t}|s_{t},h_{t})$ to verify our proposal since geometrical information remains invariant to appearance variations. The regularization task enforces the latent state to predict consistent pixel-wise depth under image augmentations by minimizing the negative of log-likelihood loss ($\mathcal{L}_{t}^{d}=-\ln q(o^{d}_{t}|s_{t},h_{t})$). We have also experimented with image denoising and inpainting on DeepMind control suite [33], where depth is not available. However, denoising and impainting are not truly invariant to RGB observation hence the performance gain on DeepMind control suite is limited. Other examples of intervention invariant auxiliary tasks for our augmentations are dense semantic segmentation, dense scene flow or sparse landmarks detection. Furthermore, our idea paves the foundation and generalizes easily to other interventions for other future tasks. The key challenge is to design different interventions and intervention invariant auxiliary tasks.
@@ 3.1.3 Learning Latent Dynamics
We leverage the DreamerV2 approach to train our world model. Similar to DreamerV2, the latent dynamics is modeled as a recurrent state space model which is relies on recurrent neural network $f_{\theta}$ that utilizes its deterministic hidden state $h_{t}$ to predict the prior stochastic state $\widehat{s}_{t}$. This enables efficient latent imagination for planning [10, 12]. In summary, dynamics models and representation learning modules are tightly integrated as a world model and have the following components,
@@
The world model is represented by neural networks and $\theta$ represents their combined parameters. It is optimized jointly with invariant representation learning module and intervention-invariant regularizer, by minimizing following loss,
@@
where, $\mathcal{L}_{t}^{d}=-\ln q(o^{d}_{t}|s_{t})$, $\mathcal{L}_{t}^{d}=-\ln q(r_{t}|s_{t})$ and $\mathcal{L}^{KL}_{t}=KL(p(s_{t}|s_{t-1},a_{t-1},\tilde{s_{t}})||q(s_{t}|s_{t-1},a_{t-1}))$.
@@ 3.1.4 Learning Controller
The objective of the controller is to optimize the expected rewards of the action, which is optimized using an actor critic approach. The actor critic approach considers the rewards beyond the horizon. Inspired by world model [10] and Dreamer [12], we learn an action model and a value model in the imagined latent space of the world model. The action model implements a policy that aims to predict future actions that maximizes the total expected rewards in the imagined environment. Given $H$ as the imagination horizon length, $\gamma$ the discount factor for the future rewards, action and policy model are defined as follows:
@@ 3.2 Implementation Details
The proposed ReCoRe expands on the publicly available code base of DreamerV2 [13]. Following MoCo [16] and BYOL [8] we have used the moving average version of the query encoder to encode the keys $K$ with a momentum value of 0.999. The contrastive loss is jointly optimized with the world model using Adam [17]. We have used five layer encoder with starting number of feature maps equal to 32, then doubled in every consecutive layers. To encode the task observations we used two dense layers of size 32 with ELU activations [5]. The features from RGB image observation and task observation are concatenated before sending to the representation model. Replay buffer capacity is $3e^{5}$ for both 100k and 500k steps experiments. We update the model parameters on every fifth interactive steps. Further, all architectural details and hyperparameters are provided in the appendix. The training time of ReCoRe is around 3 days on a workstation with two Nvidia GeForce RTX 3090 for 500k steps, which is twice higher than closest state-of-the-art model-based RL model DreamerV2 [13].
@@
Table 1: Out-of-distribution generalization results on PointGoal navigation task from iGibson 1.0 dataset. Success rate (SR) and Success weighted by (normalized inverse) Path Length (SPL) are shown. We have trained on five scenes, and tested on held-out three scenes and visual textures. Our proposed world model with ReCoRe outperforms state-of-the-art RL models RAD, CURL, DreamerV2 and MWM and DV2+G DINO on 500K interactive steps, while it is on par with DV2+G DINO on 100K steps. Even though data augmentation (DA) improves the DreamerV2, proposed invariant features learning technique with contrastive loss (CL) and intervention invariant auxiliary task (D) is significantly better. ReCoRe collapses when we remove the D, however replacing with common RGB image (I) reconstruction recovers the performance only slightly. Where, -- denotes ‘without’ and $+$ denotes ‘with’.
@@ 4 Experiments
In our evaluation we test ReCoRe on 3 datasets and compare to the state of the art in model-based and model-free RL. Specifically, we use the PointGoal navigation task from the iGibson 1.0 environment [30] to evaluate out-of-distribution (OoD) generalization, and include the Gibson dataset [37] to test sim-to-real performance. Here we follow common practice as we compare performance based on Success Rate (SR) and Success weighted by (normalized inverse) Path Length (SPL) at 100k and 500k environment steps, which tests sample efficiency [20, 19, 38, 13]. We further report on results for the DMControl suite [33], and present an ablation study.
@@ 4.1 Baselines
We compare our approach against state-of-the-art model-free RL methods RAD [19] and CURL [20], model-based RL method DreamerV2 [13], recent transformer based Masked World Model (MWM) [29] and language guided foundation model Grounding DINO [23]. We train RAD, CURL, DreamerV2 and MWM from scratch with the respective hyperparameters proposed by the authors using the official source code provided. For Grounding DINO we freeze the visual encoder and train world-model and policy (DreamerV2) from scratch using the features from the visual encoder, which we refer to as DV2+G DINO. We believe that DV2+G DINO is the strongest baseline since it has been pretrained on large amount of data, furthermore language guidance makes the representation higly robust. Though our experiments in Table 1,2 and 3, it can be seen that (i) RAD simply overfits to the in-distribution data, thus after observing large amount data (500K environment steps) its performance improves over CURL. (ii) DreamerV2 lacks any data augmentation in the encoder like CURL/RAD, therefore it performs significantly worse for OoD generalization. (iii) MWM relies on masking that forces the encoder to learn missing data, therefore it performs slightly better than DreamerV2. However, masking only helps in interpolation and not in extrapolation, consequently the OoD generalization is poor. (iv) Being a pretrianed model G DINO shows high sample efficiency for smaller amount of data and is on par with our approach. However, after observing sufficient amount of data our approach shows significantly better performance.
@@
Table 2: iGibson-to-Gibson dataset: sim-to-real perception transfer results on navigation task. We choose success rate (SR) and Success weighted by (normalized inverse) Path Length (SPL) for evaluation of the models. We have trained the models on the artist created textures of iGibson, and tested on held-out five scenes from the Gibson, which are 3D scan of the real scenes. Our proposed ReCoRe outperforms state-of-the-art RL models RAD, CURL and even madel-based RL combined with language guided foundation models on 100k and 500k interactive steps.
@@ 4.2 Out-of-Distribution Generalization
We have tested our proposed ReCoRe on random PointGoal navigation tasks of the iGibson 1.0 environment [30] for OoD generalization. The iGibson dataset contains 15 floor scenes with 108 rooms. The scenes are replicas of real-world homes with artist designed textures and materials. RGB, depth and task related observation (i.e. goal location, current location, and linear and angular velocities of the robot) are used. We emphasize, depth is only used during training. Actions includes rotation in radians and forward distance in meters for the TurtleBot. We split iGibson for OoD generalization, as we choose five scenes for training and tested on the held-out three scenes. We also held-out visual textures for all object classes. The details are provided in the appendix.
100k Steps Total Rewards ReCoRe CURL Dreamer RAD SAC+AE Pixel SAC State SAC
Finger, spin 486$\pm$191 767$\pm$56 341$\pm$70 856$\pm$73 740$\pm$64 179$\pm$66 811$\pm$46
Cartpole, swingup 472$\pm$67 582$\pm$146 326$\pm$27 828$\pm$27 311$\pm$11 419$\pm$40 835$\pm$22
Reacher, easy 327$\pm$98 538$\pm$233 314$\pm$155 826$\pm$219 274$\pm$14 145$\pm$30 746$\pm$25
Cheetah, run 321$\pm$78 299 $\pm$48 235$\pm$ 137 447$\pm$88 267$\pm$24 197$\pm$15 616$\pm$18
Walker, walk 654$\pm$100 403$\pm$24 277$\pm$12 504$\pm$191 394$\pm$22 42$\pm$12 891$\pm$82
Ball in cup, catch 830$\pm$118 769$\pm$43 246$\pm$174 840$\pm$179 391$\pm$82 312$\pm$63 746$\pm$91
@@
Finger, spin 471$\pm$173 926$\pm$45 796$\pm$183 947$\pm$101 884$\pm$128 179$\pm$166 923$\pm$21
Cartpole, swingup 675$\pm$64 841$\pm$45 762$\pm$27 863$\pm$9 735$\pm$63 419$\pm$40 848$\pm$15
Reacher, easy 891$\pm$72 929$\pm$44 793$\pm$164 955$\pm$71 627$\pm$58 145$\pm$30 923$\pm$24
Cheetah, run 633$\pm$70 518$\pm$28 570$\pm$253 728$\pm$71 550$\pm$34 197$\pm$15 795$\pm$30
Walker, walk 965$\pm$4 902$\pm$43 897$\pm$49 918$\pm$16 847$\pm$48 42$\pm$12 948$\pm$54
Ball in cup, catch 950$\pm$20 959$\pm$27 879$\pm$ 87 974$\pm$12 794$\pm$ 58 312$\pm$ 63 974$\pm$33
Table 3: Experiment results on DMControl. Results are reported as averages across 10 seeds. ReCoRe achieves state-of-the-art performance on the 5 out of 12 experiments, where models are targeted for other (new) downstream tasks as well. Additionally we report RAD [19], SAC+AE [39], Pixel SAC [11] and State SAC [11] which do not consider additional downstream tasks. However, as explained in the literature review if an agent does not need to consider the downstream tasks then optimization of the additional constraints as in ReCoRe and CURL is not necessary, and thus performance is expected to improve (this is also noted by CURL [20]). Thus ReCoRe, CURL and Dreamer form the main comparison, as these methods optimize the features learning and the controller separately.
We report the average SR and SPL on the held-out data in the Table 1 after three training runs with random seeds. Our proposed ReCoRe outperforms state-of-the-art model-based RL method DreamerV2 and model-free methods RAD and CURL on 100k and 500k interactive steps. Since DreamerV2 is natively not trained with data augmentation (DA), we include DreamerV2 + DA in our evaluation to show a fair comparison as all other methods contain DA inherently. We furthermore, include transformer backed MWM and language guided foundation model DV2+ G DINO Nevertheless, the proposed ReCoRe still outperforms the competitors through the invariant features learning technique with contrastive loss (CL) and regularization in the form of invariant auxiliary task (D). Further, our results show that data augmentation can also improve model-based RL, which was previously shown only for model-free RL methods [19].
@@ 4.3 Sim-to-Real Transfer
We use the Gibson dataset [37] for sim-to-real transfer experiments of the perception module, i.e. representation learning module of the world model; however please note that the robot controller is still a part of the simulator. Gibson scenes are created by 3D scanning of the real scenes, and it uses a neural network to fill the pathological geometric and occlusion errors. We have trained all models on the artist created textures of iGibson and tested on five scenes from Gibson. The results are shown in Table 2. Our proposed ReCoRe significantly outperforms RAD, CURL and is slightly better compared to DV2+G DINO on 100k interactive steps, while on 500k steps ReCoRe surpasses all the baselines with significant margin. This shows that ReCoRe learns more stable features and is better suited for sim-to-real transfer.
@@ 4.4 Generalizable Auxiliary Tasks
The results for the DMControl suite [33] experiments are shown in Table 3. We have used image denoising and region filling as the intervention invariant auxiliary tasks in this experiments, where interventions are adding color noise and image sub-region masking. ReCoRe achieved competitive results, and the key findings are: i) even though depth reconstruction is an ideal task to enforce the invariant features learning on ReCoRe explicitly, the competitive results, even without depth reconstruction show the wider applicability of the proposed model; ii) ReCoRe is competitive with CURL, the closest state-of-the-art RL method with contrastive learning. Hence, we can conclude that ReCoRe is also competitive with end-to-end deep RL techniques even when training and evaluation environments come from similar distributions (and no OoD generalization is necessary).
Additionally we include other baseline models RAD [19], SAC+AE [39], Pixel SAC [11] and State SAC [11]. However, as explained in the literature review if an agent does not need to consider the downstream tasks then optimization of the additional constraints (as in ReCoRe and CURL) hinders the performance (also noted by CURL [20]). So ReCoRe, CURL and Dreamer are the main competitors, as they optimize the feature learning and the controller separately.

### v2

@@ 3.1.2 Intervention-invariant Regularizer
Model-based RL approaches such as DreamerV2, separate the training of the world model and controller. This makes model-based RL more sample efficient as compared to model-free RL, since future predictions from latent dynamics (rollouts) can be used to learn RL policy. However, the policy learning is disconnected from world-model learning. Therefore, the supervisory signal from the policy does not optimize the encoder. While the image reconstruction loss in DreamerV2 caters to this issue for in-distribution learning, it is not enough to overcome the out-of-distribution generalization. Contrastive learning further elevates the issue, since it trains the encoder under different data augmentations while the world model still aims to predict the unaugmented image using latent dynamics. This leads to the collapse of encoder features due to a lack of correct supervisory signals. Therefore the brute-force combination of world-model and contrastive loss fails (as can be seen in Table 1, ReCoRe vs. ReCoRe-D). To this end, we propose regularization in the form of an intervention invariant auxiliary task to explicitly enforce invariance, which is robust against changes to the nuisance variables. For navigation tasks, we choose depth reconstruction $p_{\theta}(\widehat{d}_{t}|s_{t},h_{t})$ to verify our proposal since geometrical information remains invariant to appearance variations. The regularization task enforces the latent state to predict consistent pixel-wise depth under image augmentations by minimizing the negative of log-likelihood loss ($\mathcal{L}_{t}^{d}=-\ln q(o^{d}_{t}|s_{t},h_{t})$). We also experiment with image denoising and semantic segmentation on DeepMind control suite [33], where depth is not available. However, denoising is not truly invariant to RGB observation hence the performance gain on DeepMind control suite is limited but dense semantic segmentation improves the results. Other examples of intervention invariant auxiliary tasks for our augmentations are dense scene flow or sparse landmarks detection. Furthermore, our idea paves the foundation and generalizes easily to other interventions for future tasks. The key challenge is to design different interventions and intervention-invariant auxiliary tasks.
@@ 3.1.3 Learning Latent Dynamics
We leverage the DreamerV2 approach to train our world model. Similar to [13], the latent dynamics are modeled as a recurrent state space model which relies on recurrent neural network $f_{\theta}$ that utilizes its deterministic hidden state $h_{t}$ to predict the prior stochastic state $\widehat{s}_{t}$. This enables efficient latent imagination for planning [10, 12]. Thus, dynamics models and representation learning modules are tightly integrated as world model and have the following components:
@@
The world model is represented by neural networks and $\theta$ represents their combined parameters. It is optimized jointly with the invariant representation learning module and intervention-invariant regularizer, by minimizing,
@@
where, $\mathcal{L}_{t}^{d}=-\ln q(o^{d}_{t}|s_{t})$, $\mathcal{L}_{t}^{r}=-\ln q(r_{t}|s_{t})$ and $\mathcal{L}^{KL}_{t}=KL(p(s_{t}|s_{t-1},a_{t-1},\tilde{s_{t}})||q(s_{t}|s_{t-1},a_{t-1}))$.
@@ 3.1.4 Learning Controller
The objective of the controller is to optimize the expected rewards of the action, which is optimized using an actor critic approach. The actor critic approach considers the rewards beyond the horizon. Inspired by world model [10] and Dreamer [12], we learn an action model and a value model in the imagined latent space of the world model. The action model implements a policy that aims to predict future actions that maximizes the total expected rewards in the imagined environment. Given $H$ as the imagination horizon length and $\gamma$ as the discount factor for the future rewards, the action and value model are defined as follows:
@@ 3.2 Implementation Details
The proposed ReCoRe expands on the publicly available code base of DreamerV2 [13]. Following MoCo [16] and BYOL [8] we have used the moving average version of the query encoder to encode the keys $K$ with a momentum value of 0.999. The contrastive loss is jointly optimized with the world model using Adam [17]. We have used five layers encoder with a starting number of feature maps equal to 32, then doubled in every consecutive layer. To encode the task observations we used two dense layers of size 32 with ELU activations [5]. The features from RGB image observation and task observation are concatenated before sending to the representation model. Replay buffer capacity is $3e^{5}$ for both 100k and 500k steps experiments. We update the model parameters on every fifth interactive step. Further, all architectural details and hyperparameters are provided in the appendix. The training time of ReCoRe is around 3 days on a workstation with two Nvidia GeForce RTX 3090 for 500k steps, which is twice higher than the closest state-of-the-art model-based RL model DreamerV2 [13].
@@
Table 1: Out-of-distribution generalization results on PointGoal navigation task from iGibson 1.0 dataset. Success rate (SR) and Success weighted by (normalized inverse) Path Length (SPL) are shown. We have trained on five scenes, and tested on held-out three scenes and visual textures. Our proposed world model with ReCoRe outperforms state-of-the-art RL models RAD, CURL, DreamerV2 and MWM and DV2+G DINO on 500K interactive steps, while it is on par with DV2+G DINO on 100K steps. Even though data augmentation (DA) improves the DreamerV2, the proposed invariant features learning technique with contrastive loss (CL) and intervention invariant auxiliary task (D) is significantly better. ReCoRe collapses when we remove the auxiliary D, however replacing with common RGB image reconstruction (I) recovers the performance slightly. Where, -- denotes ‘without’ and $+$ denotes ‘with’.
@@ 4 Experiments
In our evaluation, we test ReCoRe on 3 datasets and compare it to the state of the art in model-based and model-free RL. Specifically, we use the PointGoal navigation task from the iGibson 1.0 environment [30] to evaluate out-of-distribution (OoD) generalization, and include the Gibson dataset [37] to test sim-to-real performance. Here we follow common practice as we compare performance based on Success Rate (SR) and Success weighted by (normalized inverse) Path Length (SPL) at 100k and 500k environment steps, which tests sample efficiency [20, 19, 38, 13]. We further report on results for the DMControl suite [33], and present an ablation study.
@@ 4.1 Baselines
We compare our approach against state-of-the-art model-free RL methods RAD [19] and CURL [20], model-based RL method DreamerV2 [13], recent transformer based Masked World Model (MWM) [29] and language guided foundation model Grounding DINO [23]. We train RAD, CURL, DreamerV2 and MWM from scratch with the respective hyperparameters proposed by the authors using the official source code provided. For Grounding DINO we freeze the visual encoder and train the world model and policy (DreamerV2) from scratch using the features from the visual encoder, which we refer to as DV2+G DINO. We believe that DV2+G DINO is the strongest baseline since it has been pretrained on a large amount of data, furthermore language guidance makes the representation highly robust. Through our experiments in Table 1,2 and 3, it can be seen that (i) RAD simply overfits to the in-distribution data, though after observing large amounts of data (500K environment steps) its performance improves over CURL. (ii) DreamerV2 lacks any data augmentation in the encoder like CURL/RAD, therefore it performs significantly worse for OoD generalization. (iii) MWM relies on masking that forces the encoder to learn missing data, therefore it performs slightly better than DreamerV2. However, masking only helps in interpolation and not in extrapolation, consequently the OoD generalization is poor. (iv) Being a pretrained model DV2+G DINO shows high sample efficiency for a smaller amount of data and is on par with our approach. However, after observing a sufficient amount of data our approach shows significantly better performance.
@@
Table 2: iGibson-to-Gibson dataset: sim-to-real perception transfer results on navigation task. We choose success rate (SR) and Success weighted by (normalized inverse) Path Length (SPL) for evaluation of the models. We have trained the models on the artist created textures of iGibson, and tested on five held-out scenes from the Gibson Dataset, which are 3D scan of the real scenes. Our proposed ReCoRe outperforms state-of-the-art RL models RAD, CURL and even model-based RL combined with language guided foundation models (DV2+G DINO) on 100k and 500k interactive steps.
@@ 4.2 Out-of-Distribution Generalization
We have tested our proposed ReCoRe on random PointGoal navigation tasks of the iGibson 1.0 environment [30] for OoD generalization. The iGibson dataset contains 15 floor scenes with 108 rooms. The scenes are replicas of real-world homes with artist designed textures and materials. RGB, depth and task related observation (i.e. goal location, current location, and linear and angular velocities of the robot) are used. We emphasize, depth is only used during training. Actions include rotation in radians and forward distance in meters for the TurtleBot. We split iGibson for OoD generalization, as we chose five scenes for training and tested on the held-out three scenes. We also held out visual textures for all object classes. The details are provided in the appendix.
100k Steps Total Rewards ReCoRe+I ReCoRe+S CURL Dreamer RAD SAC+AE Pixel SAC
Finger, spin 486$\pm$191 474$\pm$53 767$\pm$56 341$\pm$70 856$\pm$73 740$\pm$64 179$\pm$66
Cartpole, swingup 472$\pm$67 449$\pm$121 582$\pm$146 326$\pm$27 828$\pm$27 311$\pm$11 419$\pm$40
Reacher, easy 327$\pm$98 982$\pm$9 538$\pm$233 314$\pm$155 826$\pm$219 274$\pm$14 145$\pm$30
Cheetah, run 321$\pm$78 400$\pm$56 299 $\pm$48 235$\pm$ 137 447$\pm$88 267$\pm$24 197$\pm$15
Walker, walk 654$\pm$100 739$\pm$133 403$\pm$24 277$\pm$12 504$\pm$191 394$\pm$22 42$\pm$12
Ball in cup, catch 830$\pm$118 859$\pm$287 769$\pm$43 246$\pm$174 840$\pm$179 391$\pm$82 312$\pm$63
@@
Finger, spin 471$\pm$173 591$\pm$181 926$\pm$45 796$\pm$183 947$\pm$101 884$\pm$128 179$\pm$166
Cartpole, swingup 675$\pm$64 777$\pm$64 841$\pm$45 762$\pm$27 863$\pm$9 735$\pm$63 419$\pm$40
Reacher, easy 891$\pm$72 955$\pm$38 929$\pm$44 793$\pm$164 955$\pm$71 627$\pm$58 145$\pm$30
Cheetah, run 633$\pm$70 731$\pm$51 518$\pm$28 570$\pm$253 728$\pm$71 550$\pm$34 197$\pm$15
Walker, walk 965$\pm$4 960$\pm$2 902$\pm$43 897$\pm$49 918$\pm$16 847$\pm$48 42$\pm$12
Ball in cup, catch 950$\pm$20 984$\pm$5 959$\pm$27 879$\pm$ 87 974$\pm$12 794$\pm$ 58 312$\pm$ 63
Table 3: Experiment results on DMControl. Results are reported as averages across 10 seeds. ReCoRe with segmentation (S) as a regularizer achieves state-of-the-art performance over competitors in the literature on 8 out of 12 experiments, where models are targeted for other (new) downstream tasks as well. Additionally, we report RAD [19], SAC+AE [39], and Pixel SAC [11] which do not consider additional downstream tasks. However, as explained in the literature review if an agent does not need to consider the downstream tasks then optimization of the additional constraints as in ReCoRe and CURL is not necessary, and thus performance is expected to improve (this is also noted by CURL [20]). Thus ReCoRe, CURL and Dreamer form the main comparison, as these methods optimize the features learning and the controller separately.
We report the average SR and SPL on the held-out data in Table 1 after three training runs with random seeds. Our proposed ReCoRe outperforms state-of-the-art model-based RL method DreamerV2 and model-free methods RAD and CURL on 100k and 500k interactive steps. Since DreamerV2 is natively not trained with data augmentation (DA), we include DreamerV2 + DA in our evaluation to show a fair comparison as all other methods contain DA inherently. We furthermore include transformer backed MWM and language guided foundation model DV2  + G DINO. Nevertheless, the proposed ReCoRe still outperforms the competitors through the invariant features learning technique with contrastive loss (CL) and regularization in the form of invariant auxiliary task (D). Further, our results show that data augmentation can also improve model-based RL, which was previously shown only for model-free RL methods [19].
@@ 4.3 Sim-to-Real Transfer
We use the Gibson dataset [37] for sim-to-real transfer experiments of the perception module, i.e. representation learning module of the world model; however please note that the robot controller is still a part of the simulator. Gibson scenes are created by 3D scanning real scenes, and a neural network is used to fill pathological geometric and occlusion errors. We have trained all models on the artist created textures of iGibson and tested on five scenes from Gibson. Table 2 shows the results. Our proposed ReCoRe significantly outperforms RAD, CURL and is slightly better compared to DV2+G DINO on 100k interactive steps, while on 500k steps ReCoRe significantly surpasses all the baselines. This shows that ReCoRe learns more stable features and is better suited for sim-to-real transfer.
@@ 4.4 Generalizable Auxiliary Tasks
The results for the DMControl suite [33] experiments are shown in Table 3. We have used image denoising (I) and semantic segmentation (S) as the intervention invariant auxiliary tasks in these experiments. ReCoRe achieved competitive results, and the key findings are: i) even though depth reconstruction is an ideal task to enforce the invariant features learning on ReCoRe explicitly, the competitive results, even without depth reconstruction show the wider applicability of the proposed model; ii) ReCoRe outperform CURL on 8 out of 12 experiments, the closest state-of-the-art RL method with contrastive learning; iii) better invariant regularizer yields better results i.e. ReCoRe produces better results with segmentation than denoising as an auxiliary task. Hence, we can conclude that ReCoRe is also competitive with end-to-end deep RL techniques even when training and evaluation environments come from similar distributions (and no OoD generalization is necessary).
Additionally we include other baseline models RAD [19], SAC+AE [39], and Pixel SAC [11]. However, as explained in the literature review, if an agent does not need to consider the downstream tasks then optimization of the additional constraints (as in ReCoRe and CURL) hinders the performance (also noted by CURL [20]). So ReCoRe, CURL and Dreamer are the main competitors, as they optimize the feature learning and the controller separately. We have noticed that hyperparameter optimization for individual DMC tasks yields better results. However, we used the same set of parameters for all the tasks with ReCoRe.